Grade 6Algebra

One side of a parallelogram is 10 cm and its diagonals are 16 cm and 24 cm respectively. Find its area

ryan lester abang
6 Years agoGrade 6

1 Answer

Arun
6 Years ago

The parallelogram is made up of four triangles made up of the diagonals and four sides. The areas can be found using Heron's Formula. For a triangle with three sides, a, b, & c:

s = (a+b+c) / 2; Area = squareroot(s(s-a)(s-b)(s-c))

The point of intersection of the diagonals is at the midpoint of each diagonal, since diagonals of a parallelogram bisect each other. If the three sides of one triangle is a = 8, b = 12, c = 10

s = (8+12+10) = 15

A = sqrt(15(15-8)(15-12)(15-10)) = 39.66.

The areas of the four interior triangles of a parallelogram are equal, therefore

A(parallelogram) = 4(39.7) = 159